# Human to Document, AI to Code: Three Case Studies of Comparing GenAI for Notebook Competitions

### v1

@@
Abstract.
Computational notebooks have become the preferred tool of choice for data scientists and practitioners to perform analyses and share results. Notebooks uniquely combine scripts with documentation. With the emergence of generative AI (GenAI) technologies, it is increasingly important, especially in competitive settings, to distinguish the characteristics of human-written versus GenAI.
@@ Abstract.
In this study, we present three case studies to explore potential strengths of both humans and GenAI through the coding and documenting activities in notebooks. We first characterize differences between 25 code and documentation features in human-written, medal-winning Kaggle notebooks. We find that gold medalists are primarily distinguished by longer and more detailed documentation. Second, we analyze the distinctions between human-written and GenAI notebooks. Our results show that while GenAI notebooks tend to achieve higher code quality (as measured by metrics like code smells and technical debt), human-written notebooks display greater structural diversity, complexity, and innovative approaches to problem-solving. Based on these results, we envision the work as groundwork that highlight four agendas to further investigate how GenAI could be utilized in notebooks that maximizes the potential collaboration between human and AI.
Empirical Study, Notebooks, GenAI Code
conference: AIware 2025; November 19–20, 2025; Seoul, South Korea
1. Introduction
@@
Computational notebooks have rapidly become an important tool for data scientists and researchers. According to Jupyter, a notebook is a shareable document that combines code, text, data, and rich visualizations, offering an interactive environment for prototyping, data exploration, and sharing ideas, so that users can learn coding and documentation [^1] https://docs.jupyter.org/en/latest/. Platforms like Kaggle have further accelerated the adoption of notebooks, offering a space for machine learning competitions that attract both industry practitioners and academic researchers (Perkel, 2018). As Kaggle [^2] https://www.kaggle.com/ competitions become more popular and prestigious, participants face increasingly tougher competition to create the best, most insightful notebooks.
@@
The landscape of coding has been transformed by the emergence of generative AI (GenAI) technologies such as ChatGPT [^3] https://chatgpt.com/, Gemini [^4] https://gemini.google.com/ Claude [^5] https://claude.ai/ and open models from Meta Ollama [^6] https://ollama.com/ and Huggingface [^7] https://huggingface.co/. These tools have the potential to assist practitioners by generating code, writing explanations, and even producing entire notebooks with minimal human intervention. However, this technological shift raises important questions: How can we distinguish between notebooks created by humans and those generated by GenAI? More importantly, what can we learn from each to advance the state of the art in computational notebooks? While prior work has examined code quality (Grotov et al., 2022), reproducibility (Pimentel et al., 2021), and collaboration (Wang et al., 2019) in notebooks, the specific contributions and limitations of GenAI notebooks remain underexplored. Competitions provide an ideal experimental setting, as participants are incentivized to submit their highest-quality work for rigorous community assessment and ranking.
@@
Figure 1. Overview of Data Collection:
RQ1: What are differences between documentation and code for human-written notebook? The motivation behind RQ1 is to characterize the fundamental code and documentation patterns in human-written notebooks, revealing essential programming behaviors, documentation styles, and structural approaches that differentiate gold-medal human notebooks from the rest, establishing a crucial baseline.
RQ2: What are the key differences between GenAI and human-written data science notebooks? The motivation behind RQ2 is to identify the distinguishing factors that separate human and GenAI approaches in competitive data science, focusing on the impact of code quality metrics versus documentation on medal-worthy outcomes.
Our findings show that GenAI notebooks achieve higher code quality, with significantly fewer code smells, technical debt, and coding violations. However, human-written notebooks stand out for their more comprehensive and accessible documentation. In particular, gold medal-winning notebooks feature more than twice as much markdown content and narrative explanation as non-gold group. Moreover, while GenAI documentation tends to require a higher reading level, human-written explanations are generally clearer and more approachable.
2. Three Case Studies of Notebook Competitions
@@
3. Study Design
Figure 1 presents the workflow of the data collection process. We utilized the KGTorrent(Quaranta et al., 2021) tool to retrieve Jupyter notebooks. KGTorrent constitutes a comprehensive dataset encompassing computational notebooks (code kernels) and their associated metadata sourced from the Kaggle platform.
To obtain an updated version of the KGTorrent dataset, Kaggle metadata [^12] https://www.kaggle.com/datasets/kaggle/meta-kaggle was downloaded on May 25th, 2025, and Jupyter notebooks were subsequently retrieved through API requests using the KGTorrent tool. This process yielded a total of 205,269 notebooks. The data collection process consists of four main stages: first, we identify target competitions for analysis; second, we apply data filtering criteria to select relevant notebooks; third, we generate corresponding GenAI notebooks using GenAI models; and finally, we extract features from both human-written and GenAI notebooks for comparative analysis. To facilitate replication studies, our scripts and dataset are publicly available online, which can be found on
3.1. Data Collection and Filtering
The filtering process yielded a total of 465 notebooks, with the distribution of medal achievements across the three competitions presented in Table 1. In this study, the notebook medals (i.e., Bronze, Silver, Gold) serve as proxies for computational notebook quality, as medal progression and achievement represent community recognition and appreciation of the work’s merit. Following competition selection, a systematic filtering process was implemented to ensure dataset integrity and completeness using the following criteria:
  1. (1)
@@ 3.1. Data Collection and Filtering
Table 1. Medal Distribution Across Competitions
@@
3.2. Generating notebook using GenAI
– Used prompt
After .ipynb files in JSON format were generated by the GenAI models, the JSON format files were converted to standard Jupyter Notebook format and manually checked for file integrity and structural validity to ensure proper formatting.
3.3. Extracting features
Individual features were extracted resulting in a total of 25 features:
This extraction process utilized the SonarQube [^13] https://sonarcloud.io/explore/projects tool, an open-source static analysis platform designed for continuous code quality monitoring that integrates analysis and reporting throughout the software development process (Murillo and Jenkins, 2021; Campbell and Papapetrou, 2013; Lenarduzzi et al., 2020) is used to extract some features(*) through API.
Table 2. Statistical Comparisons of Gold and Non-Gold Human-Written Notebooks
@@ 3.3. Extracting features
Table 3. Gold vs Non-Gold Feature Statistics Comparisons
@@
4. Empirical Study
@@
4.1. Answering RQ1
To address RQ1, we assessed data normality through the Shapiro-Wilk test (Shapiro and Wilk, 1965) and analyzed statistically significant differences between notebook features and medal achievement using the Kruskal-Wallis H test (Kruskal and Wallis, 1952) with effect size calculations using the human-written notebooks dataset illustrated in Figure  1.
Table 2 shows that gold medalists primarily differentiate themselves through documentation practices rather than coding metrics. Documentation features demonstrated the strongest discriminating power between gold and non-gold medalists, with markdown characters, lines, and sentence counts showing highly significant differences (p-value $\leq$ 0.001) and medium effect sizes (0.056-0.085).
The distribution analyses in Table 3 confirm that gold medalists consistently produce substantially longer documentation than non-gold medalists. In contrast, code-related metrics showed weaker differentiation patterns. The IEEE-CIS competition displayed significance only in specific elements like duplicated lines and code character quantity, but with small effect sizes (0.033-0.046), indicating limited practical impact on distinguishing performance levels. These findings indicate that documentation, rather than code quality, primarily separates gold medalists from other competitors. Gold medalists systematically generate notebooks with more extensive markdown content.
@@ 4.1. Answering RQ1
4.2. Answering RQ2
Table 4. Statistical Comparisons of GenAI and Each Medal Notebooks
@@ 4.2. Answering RQ2
Table 5. Human vs GenAI Feature Statistics Comparison in IEEE-CIS Competition
Feature Source Mean Median Std Min Max
Functions Human 3.16 2.0 5.99 0 46
AI 0.67 0.0 1.15 0 2
Statements Human 121.49 91.5 112.14 10 586
AI 94.33 35.0 111.54 25 223
Comment Lines Human 40.93 23.5 49.54 0 252
AI 32.67 18.0 29.84 13 67
#Code Cell Human 31.36 22.0 32.82 2 205
AI 9.00 8.0 2.65 7 12
Cyclomatic Complexity Human 14.30 9.5 17.53 0 126
AI 14.67 0.0 25.40 0 44
Gunning Fog Human 10.32 9.91 7.99 0 64.61
AI 16.07 16.13 4.49 11.55 20.53
Code Smells Human 29.25 0.0 264.35 0 3166
AI 3.00 3.0 0.00 3 3
Technical Debts Human 44.39 0.0 272.18 0 3228
AI 15.00 15.0 0.00 15 15
Violations Human 29.59 0.0 264.45 0 3167
AI 3.00 3.0 0.00 3 3
@@
To investigate distinguishing characteristics between GenAI and human-written data science notebooks, we conducted statistical analyses using the Kruskal-Wallis H test (Kruskal and Wallis, 1952) across the human-written notebooks dataset and the GenAI notebook dataset, as illustrated in Figure 1, focusing on medium ($\geq$0.06) to large ($\geq$0.14) effect sizes that indicate practically meaningful differences.
Documentation Accessibility: Table 5 reveals that human-written notebooks consistently produced more accessible documentation despite GenAI’s code quality advantages. The statistical comparison shows GenAI explanatory text in the IEEE-CIS competition consistently operated at senior-college-level complexity (Gunning Fog score: 16.07), compared to human documentation which operated at high-school-senior level (Gunning Fog score: 10.32). Statistical analysis in Table 4 confirmed these readability differences across multiple competitions, with medium effect sizes observed for both gold and silver medal comparisons in IEEE-CIS (p-value $\leq$ 0.05, $\eta^{2}$ $\geq$ 0.088), while Santander showed similar patterns (p-value = 0.017, $\eta^{2}$ = 0.124). Furthermore, GenAI notebooks showed significantly fewer comment lines compared to human-written notebooks in the Home-Credit competition (p-value = 0.035, $\eta^{2}$ = 0.122, medium effect size), indicating reduced inline documentation coverage alongside the increased complexity of explanatory text.
Code Quality Feautures: Conversely, Table 4 shows that GenAI notebooks demonstrated superior performance in technical code quality measures. In the IEEE-CIS competition, GenAI notebooks significantly outperformed human-written notebooks across several code quality indicators with large effect sizes ($\geq$ 0.191): code smells (p-value = 0.004, $\eta^{2}$ = 0.209), technical debt (p-value = 0.004, $\eta^{2}$ = 0.209), and violations (p-value = 0.006, $\eta^{2}$ = 0.191). Similarly, the Home-Credit competition showed GenAI notebooks exhibiting reduced structural complexity compared to medal-winning submissions, with large effect sizes observed for functions (GenAI mean: 4.00 vs Human mean: 5.71, p-value = 0.011, $\eta^{2}$ = 0.197) and statements (GenAI mean: 144.33 vs Human mean: 194.38, p-value = 0.025, $\eta^{2}$ = 0.143).Additionally, medium effect sizes were found for the number of code cells (GenAI mean: 13.00 vs Human mean: 33.65, p-value = 0.049, $\eta^{2}$ = 0.103) and cyclomatic complexity (GenAI mean: 16.00 vs Human mean: 18.53, p-value $\leq$ 0.050, $\eta^{2}$ $\geq$ 0.102), indicating GenAI generates more streamlined code organization with lower algorithmic complexity.
@@
RQ2 Summary Human and GenAI notebooks have statistical differences. We identify and summarize key differences below: Documentation features - GenAI notebooks are harder to read (higher reading levels), while the human written documentation is more easier to read (gunning fog) Code features - GenAI notebooks demonstrate higher code quality with fewer code smells, lower technical debt, and reduced violations. Meanwhile, human-written code is more complex, with greater structural complexity and higher function counts and statements.
5. Discussion
@@
5.1. Implications
This study provides a comparative analysis of human-written and the raw output of GenAI without human intervention data science notebooks in competitive environments for an understanding of the baseline capabilities and inherent strengths and weaknesses of current sides. Our findings reveal distinct complementary strengths between Human and GenAI approaches that suggest strategic applications in data science practice. As a proof of concept, there are differences, and we believe that a more comprehensive study will highlight more differences between the Human and AI. Although preliminary, the result from RQ2 demonstrates that GenAI notebooks achieved significantly superior code quality metrics features (code smells, technical debt, and violations). These results indicate that GenAI tools could serve as valuable automated code review and standardization tools.
@@ 5.1. Implications
  • Agenda Three. Expand from Notebooks to Software Projects Extend the study to traditional software artifacts, including source code repositories and developer documentation, to determine whether the trade-offs between code quality and documentation clarity persist in broader software engineering settings. Additional questions that could arise as secondary research topics are like how different is AI to human documentation in code comments, documentation artifacts and in discussions with other team members.
@@ 5.2. Threats to Validity
Internal Validity Three primary threats to internal validity require consideration.

### v2

# Human to Document, AI to Code: Comparing GenAI for Notebook Competitions

@@
Abstract
@@
Computational notebooks have become the preferred tool of choice for data scientists and practitioners to perform analyses and share results. Notebooks uniquely combine scripts with documentation. With the emergence of generative AI (GenAI) technologies, it is increasingly important, especially in competitive settings, to distinguish the characteristics of human-written versus GenAI. Our new idea is to explore the strengths of both humans and GenAI through the coding and documenting activities in notebooks. We first characterize differences between 25 code and documentation features in human-written, medal-winning Kaggle notebooks. We find that gold medalists are primarily distinguished by longer and more detailed documentation. Second, we analyze the distinctions between human-written and GenAI notebooks. Our results show that while GenAI notebooks tend to achieve higher code quality (as measured by metrics like code smells and technical debt), human-written notebooks display greater structural diversity, complexity, and innovative approaches to problem-solving. Based on these early results, we highlight four agendas to further investigate how GenAI could be utilized in notebooks that maximize the potential collaboration between human and GenAI tech.
I Introduction
@@
Computational notebooks have rapidly become an important tool for data scientists and researchers. According to Jupyter, a notebook is a shareable document that combines code, text, data, and rich visualizations, offering an interactive environment for prototyping, data exploration, and sharing ideas, so that users can learn coding and documentation [^1] https://docs.jupyter.org/en/latest/. Platforms like Kaggle have further accelerated the adoption of notebooks, offering a space for machine learning competitions that attract both industry practitioners and academic researchers [14]. As Kaggle [^2] https://www.kaggle.com/ competitions become more popular and prestigious, participants face increasingly tougher competition to create the best, most insightful notebooks.
@@
The landscape of coding has been transformed by the emergence of generative AI (GenAI) technologies such as ChatGPT [^3] https://chatgpt.com/, Gemini [^4] https://gemini.google.com/ Claude [^5] https://claude.ai/ and open models from Meta Ollama [^6] https://ollama.com/ and Huggingface [^7] https://huggingface.co/. These tools have the potential to assist practitioners by generating code, writing explanations, and even producing entire notebooks with minimal human intervention. However, this technological shift raises important questions: How can we distinguish between notebooks created by humans and those generated by GenAI? More importantly, what can we learn from each to advance the state of the art in computational notebooks? While prior work has examined code quality [5], reproducibility [15], and collaboration [21] in notebooks, the specific contributions and limitations of GenAI notebooks remain underexplored. Competitions provide an ideal experimental setting, as participants are incentivized to submit their highest-quality work for rigorous community assessment and ranking.
@@
Figure 1: Overview of Data Collection
RQ1: What are differences between documentation and code for human-written notebook?
 The motivation behind RQ1 is to characterize the fundamental code and documentation patterns in human-written notebooks, revealing essential programming behaviors, documentation styles, and structural approaches that differentiate gold-medal human notebooks from the rest, establishing a crucial baseline.
RQ2: What are the key differences between GenAI and human-written data science notebooks?
 The motivation behind RQ2 is to identify the distinguishing factors that separate human and GenAI approaches in competitive data science, focusing on the impact of code quality metrics versus documentation on medal-worthy outcomes.
Our findings show that GenAI notebooks achieve higher code quality, with significantly fewer code smells, technical debt, and coding violations. However, human-written notebooks stand out for their more comprehensive and accessible documentation. In particular, gold medal-winning notebooks feature more than twice as much markdown content and narrative explanation as the non-gold group. Moreover, while GenAI documentation tends to require a higher reading level, human-written explanations are generally clearer and more approachable.
II Three Case Studies of Notebook Competitions
@@
III Study Design
Figure 1 presents the workflow of the data collection process. We utilized the KGTorrent[17] tool to retrieve Jupyter notebooks. KGTorrent constitutes a comprehensive dataset encompassing computational notebooks (code kernels) and their associated metadata sourced from the Kaggle platform.
To obtain an updated version of the KGTorrent dataset, Kaggle metadata [^12] https://www.kaggle.com/datasets/kaggle/meta-kaggle was downloaded on May 25th, 2025, and Jupyter notebooks were subsequently retrieved through API requests using the KGTorrent tool. This process yielded a total of 205,269 notebooks. The data collection process consists of four main stages: first, we identify target competitions for analysis; second, we apply data filtering criteria to select relevant notebooks; third, we generate corresponding GenAI notebooks using GenAI models; and finally, we extract features from both human-written and GenAI notebooks for comparative analysis.
III-A Data Collection and Filtering
The filtering process yielded a total of 465 notebooks, with the distribution of medal achievements across the three competitions presented in Table I. In this study, the notebook medals (i.e., Bronze, Silver, Gold) serve as proxies for computational notebook quality, as medal progression and achievement represent community recognition and appreciation of the work’s merit. Following competition selection, a systematic filtering process was implemented to ensure dataset integrity and completeness using the following criteria:
  1. 1.
@@ III-A Data Collection and Filtering
TABLE I: Medal Distribution Across Competitions
@@
TABLE II: Statistical Comparisons of Gold and Non-Gold Human-Written Notebooks
@@
TABLE III: Gold vs Non-Gold Feature Statistics Comparisons
@@
III-B Generating notebook using GenAI
– Used prompt
After .ipynb files in JSON format were generated by the GenAI models, the JSON format files were converted to standard Jupyter Notebook format and manually checked for file integrity and structural validity to ensure proper formatting. As part of a post-review validation to assess practical runnability, a basic execution is performed to check on all nine GenAI notebooks.
III-C Extracting features
Individual features were extracted, resulting in a total of 25 features:
This extraction process utilized the SonarQube [^13] https://sonarcloud.io/explore/projects tool, an open-source static analysis platform designed for continuous code quality monitoring that integrates analysis and reporting throughout the software development process [13, 2, 9] and is used to extract some features(*) through API.
IV Empirical Study
@@
IV-A Answering RQ1
To address RQ1, we assessed data normality through the Shapiro-Wilk test [19] and analyzed statistically significant differences between notebook features and medal achievement using the Kruskal-Wallis H test [8] with effect size calculations using the human-written notebooks dataset illustrated in Figure  1.
Table II shows that gold medalists primarily differentiate themselves through documentation practices rather than coding metrics. Documentation features demonstrated the strongest discriminating power between gold and non-gold medalists, with markdown characters, lines, and sentence counts showing highly significant differences (p-value $\leq$ 0.001) and medium effect sizes (0.056-0.085).
The distribution analyses in Table III confirm that gold medalists consistently produce substantially longer documentation than non-gold medalists. In contrast, code-related metrics showed weaker differentiation patterns. The IEEE-CIS competition displayed significance only in specific elements like duplicated lines and code character quantity, but with small effect sizes (0.033-0.046), indicating limited practical impact on distinguishing performance levels. These findings indicate that documentation, rather than code quality, primarily separates gold medalists from other competitors. Gold medalists systematically generate notebooks with more extensive markdown content.
@@ IV-A Answering RQ1
IV-B Answering RQ2
TABLE IV: Statistical Comparisons of GenAI and Each Medal Notebooks
@@ IV-B Answering RQ2
TABLE V: Human vs GenAI Feature Statistics Comparison in IEEE-CIS Competition
   Feature    Source    Mean    Median    Min    Max
   Functions    Human    3.16    2.0    0    46
   AI    0.67    0.0    0    2
   Statements    Human    121.49    91.5    10    586
   AI    94.33    35.0    25    223
   Comment Lines    Human    40.93    23.5    0    252
   AI    32.67    18.0    13    67
   #Code Cell    Human    31.36    22.0    2    205
   AI    9.00    8.0    7    12
   Cyclomatic Complexity    Human    14.30    9.5    0    126
   AI    14.67    0.0    0    44
   Gunning Fog    Human    10.32    9.91    0    64.61
   AI    16.07    16.13    11.55    20.53
   Code Smells    Human    29.25    0.0    0    3166
   AI    3.00    3.0    3    3
   Technical Debts    Human    44.39    0.0    0    3228
   AI    15.00    15.0    15    15
   Violations    Human    29.59    0.0    0    3167
   AI    3.00    3.0    3    3
@@
To investigate distinguishing characteristics between GenAI and human-written data science notebooks, we conducted statistical analyses using the Kruskal-Wallis H test [8] across the human-written notebooks dataset and the GenAI notebook dataset, as illustrated in Figure 1, focusing on medium ($\geq$0.06) to large ($\geq$0.14) effect sizes that indicate practically meaningful differences.
Documentation Accessibility: Table V reveals that human-written notebooks consistently produced more accessible documentation despite GenAI’s code quality advantages. The statistical comparison shows GenAI explanatory text in the IEEE-CIS competition consistently operated at senior-college-level complexity (Gunning Fog score: 16.07), compared to human documentation which operated at high-school-senior level (Gunning Fog score: 10.32). Statistical analysis in Table IV confirmed these readability differences across multiple competitions, with medium effect sizes observed for both gold and silver medal comparisons in IEEE-CIS (p-value $\leq$ 0.05, $\eta^{2}$ $\geq$ 0.088), while Santander showed similar patterns (p-value = 0.017, $\eta^{2}$ = 0.124). Furthermore, GenAI notebooks showed significantly fewer comment lines compared to human-written notebooks in the Home-Credit competition (p-value = 0.035, $\eta^{2}$ = 0.122, medium effect size), indicating reduced inline documentation coverage alongside the increased complexity of explanatory text.
Code Quality Features: Conversely, Table IV shows that GenAI notebooks demonstrated superior performance in technical code quality measures. In the IEEE-CIS competition, GenAI notebooks significantly outperformed human-written notebooks across several code quality indicators with large effect sizes ($\geq$ 0.191): code smells (p-value = 0.004, $\eta^{2}$ = 0.209), technical debt (p-value = 0.004, $\eta^{2}$ = 0.209), and violations (p-value = 0.006, $\eta^{2}$ = 0.191). Similarly, the Home-Credit competition showed GenAI notebooks exhibiting reduced structural complexity compared to medal-winning submissions, with large effect sizes observed for functions (GenAI mean: 4.00 vs Human mean: 5.71, p-value = 0.011, $\eta^{2}$ = 0.197) and statements (GenAI mean: 144.33 vs Human mean: 194.38, p-value = 0.025, $\eta^{2}$ = 0.143).Additionally, medium effect sizes were found for the number of code cells (GenAI mean: 13.00 vs Human mean: 33.65, p-value = 0.049, $\eta^{2}$ = 0.103) and cyclomatic complexity (GenAI mean: 16.00 vs Human mean: 18.53, p-value $\leq$ 0.050, $\eta^{2}$ $\geq$ 0.102), indicating GenAI generates more streamlined code organization with lower algorithmic complexity.
Execution and Reproducibility: Of the nine notebooks executed, there were three that failed to complete the execution. These execution outcomes confirm that static analysis alone is insufficient, as it misses critical data-handling and logic errors that limit the notebooks’ practical utility. The llama_home-credit and llama_IEEE notebooks failed due to runtime errors from passing non-numeric columns to the model. The GPT_IEEE notebook failed because column names were not consistently unified, leading to key mismatches during data processing.
@@
RQ2 Summary Human and GenAI notebooks have statistical differences. We identify and summarize key differences below: Documentation features - GenAI notebooks are harder to read (higher reading levels), while the human written documentation is easier to read (gunning fog) Code features - GenAI notebooks demonstrate higher code quality with fewer code smells, lower technical debt, and reduced violations. Meanwhile, human-written code is more complex, with greater structural complexity and higher function counts and statements.
V Discussion
@@
V-A Implications
This study provides a comparative analysis of human-written and the raw output of GenAI without human intervention data science notebooks in competitive environments for an understanding of the baseline capabilities and inherent strengths and weaknesses of both approaches. Our findings reveal distinct complementary strengths between Human and GenAI approaches that suggest strategic applications in data science practice. As a proof of concept, there are differences, and we believe that a more comprehensive study will highlight more differences between the Human and AI. Although preliminary, the result from RQ2 demonstrates that GenAI notebooks achieved significantly superior code quality metrics features (code smells, technical debt, and violations). These results indicate that GenAI tools could serve as valuable automated code review and standardization tools.
@@ V-A Implications
  • Agenda Three. Expand from Notebooks to Software Projects Extend the study to traditional software artifacts, including source code repositories and developer documentation, to determine whether the trade-offs between code quality and documentation clarity persist in broader software engineering settings. Additional questions that could arise as secondary research topics could explore how different AI documentation is from human documentation in code comments, documentation artifacts and in discussions with other team members.
@@ VI Threats to Validity
VI-A Internal Validity
Three primary threats to internal validity require consideration.